THE ORIGIN OF THE Ar⋯I_{2} B−X CONTINUUM EXCITATION SIGNAL BELOW AND ABOVE THE I_{2}(B) DISSOCIATION LIMIT: BOUND-FREE TRANSITIONS OF THE LINEAR COMPLEX

JOSHUA P. DARR, JOHN J. GLENNON, and \underline{RICHARD A. LOOMIS}, Department of Chemistry, Washington University, One Brookings Drive, CB 1134, Saint Louis, MO 63130.

Discrete features associated with transitions of both the T-shaped and linear Ar⋯I_{2}(X , v′′=0) complexes are observed in laser-induced fluorescence and action spectroscopy experiments performed throughout the I_{2} B−X electronic region. The binding energy of the linear conformer is directly measured to be 250(2) cm^{−1}, 10 to 15 cm^{−1} greater than that of the T-shaped conformer. Continuum signals are observed in the spectra that continue to energies well above the I_{2}(B) dissociation limit. Our results indicate that the continuum signals can be attributed to transitions of the linear Ar⋯I_{2}(X,v′′=0) complex to the inner, repulsive walls of many Ar + I_{2}(B,v′) intermolecular potentials. The excited state complexes are thus very short lived and undergo direct dissociation into Ar + I_{2}(B,v′) products. There is no evidence for the previously proposed one-atom caging mechanism.